# version 0.9

### v5

@@ 5.5 Bibliographic Remarks
Although developed in the linear setting, adaptations of the LQR controller are also widely used to control non-linear systems [Li05, LK13].
@@ 6.1.1 Online Gradient Descent
$\|\mathbf{y}_{t+1}-\mathbf{x}^{\star}\|^{2}$$\leq$ $\ \|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}+\eta_{t}^{2}\|\nabla_{t}\|^{2}-2\eta_{t}\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$
$2\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$$\leq$ $\ \frac{\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}-\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}.$ (6.4)
@@ 6.2 Regret for Control
Since this holds for any fixed disturbance sequence $\mathbf{w}_{1:T}$, it also holds for the maximum over all disturbance sequences $\mathbf{w}_{1:T}$ with $||\mathbf{w}_{t}||\leq 1$. ∎
@@ 6.3.2 Disturbance Action Controllers
$=(A+BK)\mathbf{x}_{t}+B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}$ using $\text{using }M_{0}=I$
@@
[…]
@@ 6.4 Bibliographic Remarks
Regret minimization, online learning, and online convex optimization have a rich history and have been studied for decades in the machine learning and game theory literature. The reader is referred to the comprehensive texts of [CBL06, Haz16] for a detailed treatment. Of particular
@@ 6.5 Exercises
Chapter 7 Online Nonstochatic Control
@@ 7.3 The Gradient Perturbation Controller
4:     Use control $\mbox{Use control }\mathbf{u}_{t}={K}_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}{M}_{i}^{t}\mathbf{w}_{t-i}.$
@@ 8.3 Nonstochastic System Identification
  Define $\widehat{C_{0}}=(\widehat{B},\dots\widehat{A^{k-1}B})$, $\widehat{C_{1}}=(\widehat{AB},\dots\widehat{A^{k}B})$, and return $\widehat{A},\widehat{B}$ as
@@
and for $T_{0}=\Omega\left(\frac{d^{2}k^{2}\kappa^{2}\gamma^{6}}{\varepsilon^{2}}\log\frac{1}{\Delta}\right)$, the following holds. Algorithm 5 outputs a pair $(\widehat{A},\widehat{B})$ that satisfies, with probability $1-\Delta$,
@@ 8.4 From Indentification to Nonstochastic Control
Having recovered approximate estimates $(\widehat{A},\widehat{B})$ of the system matrices, a natural question arises: how well does the Gradient Perturbation Algorithm 4 perform on the original system $(A,B)$ when provided with only approximate estimates of the same? We briefly state a result to this extent, whose proof we leave to the explorations of an interested reader.
@@
Given estimates of the system matrices $(\widehat{A},\widehat{B})$ such that
@@ 9.3.2 Linear Dynamical Predictors
The motivation for this class is the case that the system transformations are known and the system is fully observed, namely $\mathbf{C}=I$ is the identity matrix. In this situation, the optimal prediction can be calculated using the system evolution equation $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}$. It can be seen that this prediction rule is optimal for any zero-mean noise or perturbation model. However, this optimality is only true if the state is fully observable. In the next chapter, we consider optimal predictors for zero-mean i.i.d. noise in the more general partially observed case.
@@ 9.6 Bibliographic Remarks
The class $\prod^{*}$, which is the class of optimal open-loop predictors for systems with full observation, was studied in [AHL+18]. In that paper, it was shown how to learn partially observable and even marginally stable dynamical systems whose spectral radius approaches one, as long as their transition matrix is symmetric.
@@ 10.1 Observable Systems
[…]
@@ 10.2.4 Analysis
$\Sigma_{t+1}$ =$\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}\left[(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})^{\top}\right]$
@@ 10.2.6 Linear Predictors and the Kalman Filter
Consider any partially-observed linear dynamical system $(A,B,C)$ where the Kalman filtering matrix $L$ is such that the system $A-LC$ is $\gamma$-stable. The Kalman filtering policy is $\varepsilon$-contained in the class of linear policies $\Pi^{L}_{h,k,\kappa}$, where $h=k=\gamma\log\frac{\gamma T}{\varepsilon}$ and $\kappa=\gamma^{3}$.
@@ 10.3 Bayes-Optimality under Gaussian Noise
$\begin{bmatrix}\mathbf{x}_{t}|(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})\\ \mathbf{y}_{t}|(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\hat{\mathbf{x}}_{t}\\ C\hat{\mathbf{x}}_{t}\end{bmatrix},\begin{bmatrix}\Sigma_{t}&\Sigma_{t}C^{\top}\\ C\Sigma_{t}&C\Sigma_{t}C^{\top}+\Sigma_{y}\end{bmatrix}\right).$
@@
$\begin{bmatrix}\mathbf{x}\\ \mathbf{y}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mu_{x}\\ \mu_{y}\end{bmatrix},\begin{bmatrix}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\Sigma_{22}\end{bmatrix}\right),$
@@ 10.4 Conclusion
Beyond its theoretical guarantees, the Kalman filter is widely used in practical applications ranging from autonomous systems and robotics to finance and neuroscience. However, challenges remain, particularly in scenarios with non-Gaussian noise, time-varying dynamics, or adversarial disturbances. These limitations motivate extensions such as robust filtering, non-linear filtering (e.g., the extended and unscented Kalman filters) and learning-based adaptations that integrate statistical inference with modern machine learning techniques.
@@ 10.5 Bibliographic Remarks
  • Nonlinear Filtering Extensions. The Kalman filter assumes that the dynamics is linear system and the noise is Gaussian, which limits its effectiveness in nonlinear settings. The Extended Kalman Filter (see, for example, [UA21]) addresses this issue by linearizing the dynamics at each time step via a first-order Taylor series expansion. However, linearization can introduce significant errors in highly non-linear systems. Notable extensions using more sophisticated sampling include the Unscented Kalman Filter [JU04], leading to better estimation accuracy in strongly nonlinear systems.
@@
    For highly non-linear and non-Gaussian settings, particle filters [GSS93] provide a flexible alternative. These methods approximate the posterior distribution using a set of weighted samples (particles) and employ importance sampling and resampling techniques.
@@ 10.6 Exercises
$\begin{bmatrix}\mathbf{x}\\ \mathbf{y}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mu_{x}\\ \mu_{y}\end{bmatrix},\begin{bmatrix}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\Sigma_{22}\end{bmatrix}\right),$
@@ 11.1.1 The Magic of Hankel Matrices
[…]
@@
This theorem shows that the eigenvalues of a Hankel matrix decay very rapidly. An additional illustration of this fact is given in Figure 11.1, which plots the eigenvalues of the matrix $Z_{T}$ for $T=30$ on a logarithmic scale.
Figure 11.1: Eigenvalues of Hankel matrices decrease geometrically. These are the eigenvalues of the matrix $Z_{T}$, where $T=30$, plotted on a logarithmic scale.
@@ 11.2 Spectral Predictors
$\prod_{h}^{\lambda}=\left\{\pi_{M}\ \ \left|\ \ \pi_{M}(\mathbf{u}_{1:t},\mathbf{y}_{1:t})-\pi_{M}(\mathbf{u}_{1:t-1},\mathbf{y}_{1:t-1})=M_{0}\mathbf{u}_{t}+\sum_{i=1}^{h}M_{i}\cdot\phi_{i}(\tilde{\mathbf{u}}_{t-1})\right.\right\}.$
@@
[…]
$=\beta\cdot\mu_{\alpha}^{\top}(\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top})\tilde{\mathbf{u}}_{t-1}$ basis property $\mbox{basis property }\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top}=I$
@@
$\leq\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}+\mathbf{y}_{t-1}^{\pi_{2}}\|$$\triangle\mbox{-inequality}$
@@
[…]
@@
[…]
@@
[…]
@@ 11.6 Exercises
11.1. Generalize Lemma 11.3 to all real numbers in the range $[-1,1]$, instead of $[0,1]$.
@@ 12.1.1 Generalized Linear Observation Action Controllers
Generalized observation-action policies can be seen as a special case of observation-action controllers using the technique of dimension lifting as follows. Consider any partially-observed system $(A,B,C)$ and perturbation sequence $\mathbf{w}_{t}$. Consider the lifted system $(\tilde{A},\tilde{B},\tilde{C},\tilde{D})$ given by
@@ 12.3.1 Expressivity of Disturbance Response Controllers
For any partially-observed system $(A,B,C)$ and perturbation sequence $\mathbf{w}_{t}$ producing observations $\mathbf{y}_{t}$, construct a lifted system $(\tilde{A},\tilde{B},\tilde{C})$ given by
@@
Fix any sequence of control inputs $\mathbf{u}_{t}$. Let $\mathbf{x}_{t},\mathbf{y}_{t}$ be the states and observations produced by the original partially-observed system. The equivalents for the augmented system are $[{\mathbf{x}^{\mathbf{nat}}_{t}}^{\top},\Delta_{t}^{\top}]^{\top},\tilde{\mathbf{y}}_{t}$, as we have described above. Now we can verify for all time $t$ that $\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}$ and $\mathbf{x}_{t}=\mathbf{x}^{\mathbf{nat}}_{t}+\Delta_{t}$.
@@ 13.1 The Gradient Response Controller
3:     Use Control $\mbox{Use Control }\mathbf{u}_{t}=\sum_{i=0}^{h}{M}_{i}^{t}\mathbf{y}^{\mathbf{nat}}_{t-i}$
@@ 13.2 Extension to Time Varying Systems
3:     Use Control $\mbox{Use Control }\mathbf{u}_{t}=\sum_{i=0}^{h}{M}_{i}^{t}\mathbf{y}^{\mathbf{nat}}_{t-i}$

### v6

@@ 5.5 Bibliographic Remarks
Although developed in the linear setting, adaptations of the LQR controller are also widely used to control nonlinear systems [Li05, LK13].
@@ 6.1.1 Online Gradient Descent
$\|\mathbf{y}_{t+1}-\mathbf{x}^{\star}\|^{2}\ \displaystyle\leq$ $\ \|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}+\eta_{t}^{2}\|\nabla_{t}\|^{2}-2\eta_{t}\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$
$2\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})\ \displaystyle\leq$ $\ \frac{\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}-\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}.$ (6.4)
@@ 6.2 Regret for Control
Since this holds for any fixed disturbance sequence $\mathbf{w}_{1:T}$, it also holds for the maximum over all disturbance sequences $\mathbf{w}_{1:T}$ with $\mathbf{w}_{t}||\leq 1$. ∎
@@ 6.3.2 Disturbance Action Controllers
$=(A+BK)\mathbf{x}_{t}+B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}$ $\text{using }M_{0}=I$
@@
$\leq\kappa^{2}\sum_{i=h+1}^{\infty}(1-\delta)^{i}\leq\kappa^{2}\int_{i=h}^{\infty}e^{-\delta i}di$
@@ 6.4 Bibliographic Remarks
Regret minimization, online learning, and online convex optimization have a rich history and have been studied for decades in the machine learning and game theory literature. The reader is referred to the comprehensive texts of [CBL06, Haz16] for a detailed treatment.
@@ 6.5 Exercises
Chapter 7 Online Nonstochastic Control
@@ 7.3 The Gradient Perturbation Controller
4:  $\mbox{Use control }\mathbf{u}_{t}={K}_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}{M}_{i}^{t}\mathbf{w}_{t-i}.$
@@ 8.3 Nonstochastic System Identification
[…]
@@
[…]
@@ 8.4 From Indentification to Nonstochastic Control
[…]
@@
[…]
@@ 9.3.2 Linear Dynamical Predictors
The motivation for this class is the case that the system transformations are known and the system is fully observed, namely $\mathbf{C}=I$, the identity matrix. In this situation, the optimal prediction can be calculated using the system evolution equation $\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}$. It can be seen that this prediction rule is optimal for any zero-mean noise or perturbation model. However, this optimality is only true if the state is fully observable. In the next chapter, we consider optimal predictors for zero-mean i.i.d. noise in the more general partially observed case.
@@ 9.6 Bibliographic Remarks
The class $\prod^{*}$, which is the class of optimal open-loop predictors for systems with full observation, was studied in [HSZ17]. That paper showed how to learn partially observable and even marginally stable dynamical systems whose spectral radius approaches one, as long as their transition matrix is symmetric.
@@ 10.1 Observable Systems
where $A$ and $C$ are the system matrices that define the state transition and observation model, respectively. Notice that in our example, the Kalman observability matrix is given by $\hat{K}(A,C)=\begin{bmatrix}0\ 1\\ 0\ 1\end{bmatrix}$, which is degenerate. This is not by chance; the system is observable if and only if the observability matrix has full rank. We ask the reader to provide a proof of the following result in Exercise 10.6.
@@ 10.2.4 Analysis
$\Sigma_{t+1}$ $\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}\left[(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})(\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1})^{\top}\right]$
@@ 10.2.6 Linear Predictors and the Kalman Filter
Consider any partially-observed linear dynamical system $(A,B,C)$ where the Kalman filtering matrix $L$ is such that the system ALCA-LC is $\gamma$-stable. The Kalman filtering policy is $\varepsilon$-contained in the class of linear policies $\Pi^{L}_{h,k,\kappa}$, where $h=k=\gamma\log\frac{\gamma T}{\varepsilon}$ and $\kappa=\gamma^{3}$.
@@ 10.3 Bayes-Optimality under Gaussian Noise
[…]
@@
[…]
@@ 10.4 Conclusion
Beyond its theoretical guarantees, the Kalman filter is widely used in practical applications ranging from autonomous systems and robotics to finance and neuroscience. However, challenges remain, particularly in scenarios with non-Gaussian noise, time-varying dynamics, or adversarial disturbances. These limitations motivate extensions such as robust filtering, nonlinear filtering (e.g., the extended and unscented Kalman filters) and learning-based adaptations that integrate statistical inference with modern machine learning techniques.
@@ 10.5 Bibliographic Remarks
  • Nonlinear Filtering Extensions. The Kalman filter assumes that the dynamics is linear system and the noise is Gaussian, which limits its effectiveness in nonlinear settings. The Extended Kalman Filter (see, for example, [UA21]) addresses this issue by linearizing the dynamics at each time step via a first-order Taylor series expansion. However, linearization can introduce significant errors in highly nonlinear systems. Notable extensions using more sophisticated sampling include the Unscented Kalman Filter [JU04], leading to better estimation accuracy in strongly nonlinear systems.
@@
    For highly nonlinear and non-Gaussian settings, particle filters [GSS93] provide a flexible alternative. These methods approximate the posterior distribution using a set of weighted samples (particles) and employ importance sampling and resampling techniques.
@@ 10.6 Exercises
[…]
@@ 11.1.1 The Magic of Hankel Matrices
$W_{T}$ $=\int_{0}^{1}\begin{bmatrix}1&\alpha&\alpha^{2}&&\\ \alpha&\alpha^{2}&\alpha^{3}&&\\ \alpha^{2}&\alpha^{3}&\alpha^{4}&&\\ &&&...&\\ &&&&\alpha^{2T-2}\end{bmatrix}d\alpha=\begin{bmatrix}1&\frac{1}{2}&\frac{1}{3}&&\\ \frac{1}{2}&\frac{1}{3}&\frac{1}{4}&&\\ \frac{1}{3}&\frac{1}{4}&\frac{1}{5}&&\\ &&&...&\\ &&&&\frac{1}{2T-1}\end{bmatrix}$.
@@
This theorem shows that the eigenvalues of a Hankel matrix decay very rapidly. An additional illustration of this fact is given in Figure 11.1, which plots the eigenvalues of the matrix $Z_{T}$ for $T=100$ on a logarithmic scale.
Figure 11.1: Eigenvalues of Hankel matrices decrease geometrically. These are the eigenvalues of the matrix $Z_{T}$, where $T=100$, plotted on a logarithmic scale.
@@ 11.2 Spectral Predictors
[…]
@@
$=\beta\cdot\mu_{\alpha}^{\top}\tilde{\mathbf{u}}_{t-1}$ (11.2)
$=\beta\cdot\mu_{\alpha}^{\top}(\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top})\tilde{\mathbf{u}}_{t-1}$ $\mbox{basis property }\sum_{i=1}^{T}\phi_{i}\phi_{i}^{\top}=I$
@@
$\leq\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t-1}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}+\mathbf{y}_{t-1}^{\pi_{2}}\|$ -$\triangle\mbox{-inequality}$
@@
$\leq\sqrt{T}\times O\left(\int_{h}^{\infty}e^{-\frac{i}{\log T}}di\right)$ Lemma 11.4
@@
$\int_{\alpha=0}^{1}g(\alpha)d\alpha=\int_{\alpha=0}^{1}\big(\phi_{j}^{\top}\mu_{\alpha}\big)^{2}d\alpha=\int_{\alpha=0}^{1}\phi_{j}^{\intercal}\mu_{\alpha}\mu_{\alpha}^{\intercal}\phi_{j}d\alpha=\phi_{j}^{\intercal}Z_{T}\phi_{j}=\sigma_{j}.$
@@
$g_{\max}=\max_{\alpha}\left\{(\phi_{j}^{\top}\mu_{\alpha})^{2}\right\}$ $\leq 2\cdot 6\cdot\int_{\alpha}(\phi_{j}^{\top}\mu_{\alpha})^{2}d\alpha$
$=12\cdot\phi_{j}^{\top}\left(\int_{\alpha}\mu_{\alpha}\mu_{\alpha}^{\top}d\alpha\right)\phi_{j}$
@@ 11.6 Exercises
11.1. Generalize Lemma 11.3 to all real numbers in the range [1,1][-1,1], instead of $[0,1]$.
@@ 12.1.1 Generalized Linear Observation Action Controllers
[…]
@@ 12.3.1 Expressivity of Disturbance Response Controllers
[…]
@@
Fix any sequence of control inputs $\mathbf{u}_{t}$. Let $\mathbf{x}_{t},\mathbf{y}_{t}$ be the states and observations produced by the original partially-observed system. The equivalents for the augmented system are ${\mathbf{x}^{\mathbf{nat}}_{t}}^{\top},\Delta_{t}^{\top}]^{\top},\tilde{\mathbf{y}}_{t}$, as we have described above. Now we can verify for all time $t$ that $\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}$ and $\mathbf{x}_{t}=\mathbf{x}^{\mathbf{nat}}_{t}+\Delta_{t}$.
@@ 13.1 The Gradient Response Controller
3:  $\mbox{Use Control }\mathbf{u}_{t}=\sum_{i=0}^{h}{M}_{i}^{t}\mathbf{y}^{\mathbf{nat}}_{t-i}$
@@ 13.2 Extension to Time Varying Systems
3:  $\mbox{Use Control }\mathbf{u}_{t}=\sum_{i=0}^{h}{M}_{i}^{t}\mathbf{y}^{\mathbf{nat}}_{t-i}$